OpenStudy (anonymous):

RRSP Lila and Paul are investing $5000 in their Registered Retirement Savings Plans, RRSPs, at 6% per annum, compounded quarterly. How much will each investment be worth when they reach 60, if Lila makes the deposit on her 38th birthday and Paul makes the deposit on his 48th birthday?

OpenStudy (anonymous):

I tried P as the Principle invested: 5000, i, as the interest rate per compounding period = 6/100 = 0.06 n, number of compounding periods = 4, and t, time is 60? Because I got it wrong: A = P(1 + i/n)^n(t)

OpenStudy (anonymous):

The answer in the text are: $18 534. 54, $10 217. 39

OpenStudy (dumbcow):

time is not 60. t = the number of years the 5000 has been invested Lila: t = 60-38 = 22 Paul: t = 60-48 = 12

OpenStudy (anonymous):

ohhhh.. thank you!